### 27 Nov 2023 v1

# EVCap: Retrieval-Augmented Image Captioning with External Visual–Name Memory for Open-World Comprehension

Abstract
Large language models (LLMs)-based image captioning has the capability of describing objects not explicitly observed in training data; yet novel objects occur frequently, necessitating the requirement of sustaining up-to-date object knowledge for open-world comprehension. Instead of relying on large amounts of data and scaling up network parameters, we introduce a highly effective retrieval-augmented image captioning method that prompts LLMs with object names retrieved from External Visual–name memory (EVCap). We build ever-changing object knowledge memory using objects’ visuals and names, enabling us to (i) update the memory at a minimal cost and (ii) effortlessly augment LLMs with retrieved object names utilizing a lightweight and fast-to-train model. Our model, which was trained only on the COCO dataset, can be adapted to out-domain data without additional fine-tuning or retraining. Our comprehensive experiments conducted on various benchmarks and synthetic commonsense-violating data demonstrate that EVCap, comprising solely 3.97M trainable parameters, exhibits superior performance compared to other methods of equivalent model size scale. Notably, it achieves competitive performance against specialist SOTAs with an enormous number of parameters. Our code is available at https://jiaxuan-li.github.io/EVCap.
1 Introduction
Advanced image captioning based on large language models (LLMs) [3, 9, 25, 8] has focused on the approach using big-scale models trained on ever-increasingly large-scale datasets, which is no longer viable. This is because the computational cost to train the models increases exponentially and, more importantly, updating training data is almost impossible that keep pace with the growth of novel objects in our daily lives. Sustaining ever-changing object knowledge with a reasonable cost is a pressing concern in LLMs-based models to truly unlock open-world comprehension.
Figure 1: Overall comparison of our EVCap and SOTAs. (Upper) Comparison of the number of trainable parameters, CIDEr score on COCO and NoCaps datasets. The size of each circle reflects the log number of trainable parameters. EVCap (3.97M) has less trainable parameters than others while achieving comparable results with SOTAs at scale. (Lower) Generated captions by SmallCap, BLIP-2, and our EVCap for a commonsense-violating image from the WHOOPS dataset. $\times$ and $\checkmark$ indicate incorrect and correct predictions, respectively. Incorrect objects in captions are highlighted in red, while correct ones are in blue. SmallCap and BLIP-2 give incorrect predictions for “ice skates” and “wood floor”, respectively, while our EVCap utilizes an external visual–name memory to enhance attention to objects within the image, leading to superior performance for image captioning.
Retrieval-augmented image captioning [35, 20] is emerging as an alternative since it considerably reduces training costs in both time and data while producing encouraging results. Nonetheless, with their huge datastore, it is obvious that LLMs would imitate the given texts, limiting their ability to describe open-world objects properly. For instance, SmallCap [35] considers the words “skateboard” and “wooden floor” to be a pair regardless of visual appearances containing a commonsense-violating pair of “ice skates” and “wood floor” (Fig. 1, lower). Additionally, prompting the LLMs given a lot of retrieved texts becomes cumbersome, requiring more trainable parameters. Fig. 1 (upper) shows that the CIDEr scores obtained by a lightweight SmallCap [35] with 43M trainable parameters are far away from those obtained by a heavy REVEAL [20] with 2.1B trainable parameters. Beyond that, due to the frequent occurrence of new objects, access to their sample texts is not always feasible, making the memory utilized in [35, 20] difficult to grow. We thus aim to streamline the external memory used in previous work [35, 20] by storing a sufficiently small amount of object information. And, of course, not only does the model not stereotype the example sentences, but the number of trainable parameters would be reduced drastically as a result of the causation (Fig. 1).
We follow [40, 13] to construct a key-value memory where the key is represented by object’s features, and the value corresponds to object’s name. Unlike [40, 13], which rely on object definition as the key, our method leverages the visual appearance of the object as the key because of the abundance of object images readily available on the internet. We propose an external visual–name memory tailored for ease of expansion and cost-effectiveness in upholding up-to-date object information. We present a highly effective retrieval-augmented LLMs-based image captioning method, called EVCap, that prompts frozen LLMs with object names retrieved from our proposed memory for open-world comprehension. EVCap contains a frozen image encoder ViT [14] and Q-Former [25] with trainable image query tokens for object retrieval, an attentive fusion module, a trainable linear layer for mapping between vision and language latent spaces, and a frozen LLM decoder [10] for generating captions. Specifically, the attentive fusion module feeds retrieved object names and visual features into a customized frozen Q-Former using trainable object name query tokens to implicitly reduce the presence of superfluous object names. As a result, EVCap amounts to only 3.97M trainable parameters. Once trained, the model can be adapted to new domains and large-scale data without further fine-tuning or retraining. Our contributions are as follows:
  • We provide an extensible external visual–name memory with minimal but useful object information, which enables LLMs-based models to comprehend the open world.
  • We present a remarkably lightweight and highly efficacious retrieval-augmented image captioning EVCap with 3.97M trainable parameters.
On in-/out-domain benchmarks and synthetic commonsense-violating dataset, EVCap trained solely on COCO dataset competes with other lightweight methods by a margin while being on par with other specialist SOTAs.
2 Related Work
Image captioning aims to describe the contents of a given image. It can be roughly divided into two approaches: non-LLMs-based methods and LLMs-based ones. The former approaches [22, 42, 4] typically employ a visual encoder and a language decoder in an end-to-end fashion to generate captions. However, they are incapable of describing open-world objects. The latter one leverages pre-trained large-scale vision models (CLIP [32], ViT [12]) and LLMs (GPTs [31, 7], T5 [33], LLaMA [37]) by bridging the gap between two modalities using either pre-training with large-scale data or the learned mapper or prompt techniques. LLMs-based models [29, 25, 9, 8] demonstrate advancements in image captioning challenges, allowing the capacity to describe anything as long as pre-trained vision models can recognize it. Our method belongs to the LLMs-based approaches, but instead of relying fully on the pre-trained vision model, we use object names retrieved from the external memory to augment LLMs-based image captioning.
Novel object captioning is a branch of image captioning that describes images containing objects that were not seen during training. Non-LLMs-based methods explore more objects by learning from unpaired image-sentence sources (DCC [19], NOC [39]) or relied on novel object detectors to recognize novel concepts (NBT [28], OSCAR [26] and VinVL [45]). LLMs-based methods such as ViECap [15] leverage the pre-trained CLIP [32] to obtain object entities. Nevertheless, the cut-off in training time of the pre-trained object detector or CLIP prevents it from detecting novel objects that arise quickly in reality. Unlike earlier work, we can readily update our recognition of novel concepts by adding them to external memory, ensuring that we keep any new objects from the past and even the future.
Retrieval-augmented image captioning is a recently popular approach that augments the captioning model with retrieved information for better open-world understanding. AoANet [16] uses a memory bank of image-sentence pairs and target words. SmallCap [35] employs image-to-text retrieval to obtain sampled captions from a captions datastore. RA-CM3 [44] retrieves documents from an external memory of a mixture of text and image via a dense multimodal retriever. EXTRA [34] and Re-ViLM [43] exploit the similarity of the input image and vision candidates to retrieve captions. Unlike previous methods, our external memory contains visual–name pairs to avoid redundant information in the external captions/documents. In addition, we use an attentive fusion module to mitigate the effects of irrelevant retrieved object names on caption generation.
Figure 2: Schematic of our proposed EVCap. It consists of an external visual–name memory with image embeddings and object names (upper), a frozen ViT and Q-Former equipped with trainable image query tokens, an attentive fusion module developed by a customized frozen Q-Former and trainable object name query tokens, and a frozen LLM with a trainable linear layer (lower). The ViT and Q-Former extract learned visual features from the input image, which are then used to retrieve object names from the external memory. These retrieved object names and learned visual features undergo cross-attention in the customized Q-Former, creating refined object name features. Finally, the object name features combined with visual features are fed into the LLM post a linear layer for generating captions.
3 Proposed EVCap
3.1 Idea of EVCap
We aim to build a retrieval-augmented LLMs-based image captioning model with a sufficiently small yet informative external memory. It involves two challenges: (1) constructing an expandable external memory, and (2) building an effective LLMs-based model using retrieved object names.
As discussed above, challenge (1) can be resolved by utilizing the visual appearance of objects. However, if we restrict our memory to only a visual–name pair for each object, our memory will be lacking in diversity. Therefore, we gather several images for each target object. Additionally, we keep the synthetic images in our memory to avoid the harm that synthetic images might cause to our method, as pointed out in [18]. With the capability to collect images from the internet, EVCap can be easily expanded to include novel objects from the real world effortlessly.
We base our method on a frozen pre-trained vision model and LLM with several trainable layers (Fig. 2), giving in a model that is cheap to train. To guide the LLM, we adopt a recently popular approach called prompting as in [29, 25, 11, 35, 46]. We begin by matching the learned visual features from the input image with image embeddings stored in memory, retrieving object names. We also introduce an attentive fusion module designed to implicitly remove irrelevant retrieved names. Finally, following the attentive fusion, we combine the learned visual features and object name features to form a prompt for the LLM to generate a caption, thus addressing challenge (2).
3.2 External visual–name memory
To build the external visual–name memory, we first collect image–name pairs from the external data source. After that, we encode these images into image embeddings, which serve as keys in memory, and use their names as values.
External data source. We utilize object images from LVIS dataset [17] to construct our external visual–name memory $\mathcal{M}$. Specifically, we use 1203 objects in LVIS, where we randomly select from one to ten images for each object, amounting to 8581 object images. Furthermore, as mentioned in Sec. 3.1, we also incorporate synthetic images in our memory construction. Using stable diffusion [36], we generate five additional images for each object, with a prompt of “a photo of {object name}”, resulting in a total of $M=14596$ $(8581+5\times 1203)$ images. Each object image $X^{i}$ is associated with an object name $v^{i}$. Note that many object images may share the same object name. For the sake of simplicity, we may regard each image as corresponding to a single name. In summary, we have $M$ image–name pairs $\{(X^{i},v^{i})\}^{M}_{i=1}$ for external memory construction.
External memory construction. For each image $X^{i}$, we use a frozen vision encoder $\mathcal{E}(\cdot)$ (see Sec. 3.3 for detail) to project it into 32 embeddings with the size of $1\times 768$ each: $\{\mathbf{k}^{i}_{1},\mathbf{k}^{i}_{2},\cdots,\mathbf{k}^{i}_{32}\}=\mathcal{E}(X^{i})$. We then average 32 embeddings to produce a single embedding $\mathbf{k}^{i}$ ($1\times 768$) that serves as the key (visual) in $\mathcal{M}$. The paired object name $v^{i}$ acts as its value (name). Consequently, we have the visual–name memory =$\mathcal{M}=\{(\mathbf{k}^{i},v^{i})\}^{M}_{i=1}$ which is indexed using FAISS [21], facilitating rapid searches based on similarity measures. Our memory can be expanded effortlessly by gathering additional visual–name pairs (see Sec. 5.3).
3.3 Object names retrieval
Image encoding. We feed a frozen vision encoder $\mathcal{E}$ image $X$ and image query tokens $\mathbf{T}_{\mathrm{img}}$ to produce visual features $\mathcal{Q}$. To enable the retrieval process controllable, we make image query tokens to be trainable. Thus, the image encoding process can be summarized as $\mathcal{Q}=\mathcal{E}(X,\mathbf{T}_{\mathrm{img}})$. We use the BLIP-2 pre-trained vision encoder [25], which consists of a pre-trained vision transformer ViT-g [14] outputting image features ($257\times 1408$), and a Q-Former receiving image features producing $|\mathcal{Q}|=32$ learned visual features ($1\times 768$ each). We denote $\mathcal{Q}=\{\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{32}\}$.
Retrieval. Having obtained $\mathcal{Q}$, we calculate the cosine similarity between the query $\mathbf{q}_{j}\in\mathcal{Q}$ and the key $\mathbf{k}^{i}\in\mathcal{M}$. The similarity calculation is given by $\operatorname{SIM}(\mathbf{q}_{j},\mathbf{k}^{i})=\frac{\mathbf{\mathbf{q}_{j}^{\top}}\mathbf{k}^{i}}{\|\mathbf{q}_{j}\|\|\mathbf{k}^{i}\|}$, where $i\in[1,M]$, $j\in[1,32]$. Given each $\mathbf{q}_{j}$, we select one key with the highest similarity score, resulting in 32 key–value candidates $\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}^{32}_{j=1}$.
After that, we filter out candidates with repeated object names (values), and then select the top-K values. In particular, we determine the index $j$ from the key that has the highest $\operatorname{SIM}$ score. These selected values $v_{j}^{\mathrm{best}}$ are redefined as the new notation $v_{l}$ in the retrieved top-K object names for the input image, which can be summarized as follows:
$\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}=\arg\max_{\mathbf{k}^{i}}\operatorname{SIM}\left(\mathbf{q}_{j},\mathbf{k}^{i}\right),\\ \vspace*{-0.75\baselineskip}$
$j=\arg\max_{j}\operatorname{SIM}(\mathbf{q}_{j},\mathbf{k}_{j}^{\mathrm{best}}),v_{l}\leftarrow v_{j}^{\mathrm{best}},$
where $l\in[1,\mathrm{K}]$. As a result, the retrieved top-K object names are $\{v_{l}\}^{\mathrm{K}}_{l=1}$.
3.4 Attentive fusion
Since the object names obtained from the retrieval process may be redundant, we develop an attentive fusion module to selectively distill object name features.
The retrieved object names $\{v_{l}\}^{\mathrm{K}}_{l=1}$ are concatenated together into a sequence $\mathcal{S}$, each separated by a delimiter: $\mathcal{S}=\{v_{1},\texttt{[SEP]},v_{2},\texttt{[SEP]},\cdots,\texttt{[SEP]},v_{\mathrm{K}}\}$. The sequence $\mathcal{S}$ and visual features $\mathcal{Q}$ are fed into a customized Q-Former $\mathcal{F(\cdot)}$, which is constructed from the frozen pre-trained Q-Former as we used in vision encoder $\mathcal{E}$. Nonetheless, in order to enable object names to get attention from visual features, we switch the image embedding port and the text instruction port (see [25] for architecture detail). Like in the image encoding process in Sec. 3.3, we make the object name query tokens $\mathbf{T}_{\mathrm{obj}}$ learnable during training to assist in learning object name features related to the caption. The size of $\mathbf{T}_{\mathrm{obj}}$ is $P\times 768$, where $P$ indicates the number of object name query tokens. We get the object name features $\mathcal{V}=\mathcal{F}(\mathcal{S},\mathcal{Q},\mathbf{T}_{\mathrm{obj}})$.
3.5 Caption generation
Before inputting the visual features $\mathcal{Q}$ and object name features $\mathcal{V}$ into the LLM decoder, we concatenate ($\oplus$) them and use a linear layer $\phi(\cdot)$ to project them into the input latent space of the LLM as $\phi(\mathcal{Q}\oplus\mathcal{V})$. The LLM used for caption generation in this work is the pre-trained Vicuna-13B [10], an open-source chatbot constructed from LLaMA [37]. During training and evaluation, we design a prompt in a conversational format, that is similar to [46]:
###Human: <Img><ProjFeature></Img> Describe this image in detail. ###Assistant:
in which, ProjFeature denotes the projected feature $\phi(\mathcal{Q}\oplus\mathcal{V})$ after the linear layer. In training phase, given input caption tokens $\{c_{i}\}^{L}_{i=1}$, the LLM decoder concatenates the embedded prompt $\{\mathbf{w}_{i}\}^{N}_{i=1}$ and the embedded caption tokens $\{\mathbf{c}_{i}\}^{L}_{i=1}$ as input, and predicts the caption tokens in an autoregressive fashion, while in the evaluation phase, we only need to input the embedded prompt. We train EVCap by minimizing the cross-entropy loss in an end-to-end way: $\mathcal{L}_{\theta}=-\sum_{i=1}^{L}\log p_{\theta}\left(c_{i}\mid\mathbf{w}_{1},...\mathbf{w}_{N},\mathbf{c}_{1},...,\mathbf{c}_{i-1}\right)$, in which $\theta$ indicates the trainable parameters.
4 Experimental Settings
4.1 Training setup
Implementation. EVCap uses the same image encoder as in BLIP-2 [25], consisting of a ViT-g [14] and their pre-trained Q-Former. Since we intend to obtain object name features through cross-attention between retrieved object names and visual features, we develop a customized Q-Former, which consists of BERT [23] with cross-attention layers inserted at every other transformer block. We use a frozen Vicuna-13B [10] as the caption generator.
Training dataset. For all experiments, we exclusively train EVCap using the training set of COCO dataset [27], consisting of 82k images and 5 captions per images. The entire training process takes about 3 hours on 4 A6000 GPUs, using mixed precisions (more details in the supplementary).
Table 1: Quantitative comparison against SOTA methods on three common image captioning benchmarks. * denotes using a memory bank. We report the size of training data and parameters; BLEU@4 (B@4), METEOR (M), CIDEr (C), and SPICE (S) scores on COCO test set; C and S scores on in-domain, near-domain, out-domain and overall data of NoCaps validation set; C and S scores on Flickr30k test set. Higher score is better. Bold indicates the best results among compared methods, normal indicates the second best results.
Method Training COCO NoCaps val Flickr30k
Data Para. Test In-domain Near-domain Out-domain Overall Test
B@4 M C S C S C S C S C S C S
Heavyweight-training models
VinVL [45] 8.9M 110M 38.2 30.3 129.3 23.6 96.8 13.5 90.7 13.1 87.4 11.6 90.9 12.8
AoANet+MA* [16] COCO 38.0 28.7 121.0 21.8
NOC-REK* [40] COCO 110M 104.7 14.8 100.2 14.1 100.7 13.0 100.9 14.0
RCA-NOC* [13] COCO 110M 37.4 29.6 128.4 23.1 92.2 12.9 87.8 12.6 87.5 11.5 88.3 12.4
ViECap ${}_{\text{GPT2}}$ [15] COCO 124M 27.2 24.8 92.9 18.2 61.1 10.4 64.3 9.9 65.0 8.6 66.2 9.5 47.9 13.6
InstructBLIP ${}_{\text{Vicuna-13B}}$ [11] 129M 188M 121.9 82.8
OSCAR [26] 4.1M 338M 37.4 30.7 127.8 23.5 83.4 12.0 81.6 12.0 77.6 10.6 81.1 11.7
BLIP [24] 129M 446M 40.4 136.7 114.9 15.2 112.1 14.9 115.3 14.4 113.2 14.8
BLIP-2 ${}_{\text{FlanT5-XL}}$ [25] 129M 1.2B 42.4 144.5 123.7 16.3 120.2 15.9 124.8 15.1 121.6 15.8
REVEAL* ${}_{\text{T5}}$ [20] 1.3B 2.1B 145.4 123.0
Lightweight-training models
MiniGPT4 ${}_{\text{Vicuna-13B}}$ [46] 5M 3.94M 38.0 29.6 129.6 23.4 99.0 14.8 106.9 15.3 110.8 14.9 108.8 15.1 78.4 16.9
SmallCap* ${}_{\text{GPT2}}$ [35] COCO 7M 37.0 27.9 119.7 21.3 60.6
ClipCap ${}_{\text{GPT2}}$ [29] COCO 43M 33.5 27.5 113.1 21.1 84.9 12.1 66.8 10.9 49.1 9.6 65.8 10.9
EVCap* ${}_{\text{Vicuna-13B}}$ COCO 3.97M 41.5 31.2 140.1 24.7 111.7 15.3 119.5 15.6 116.5 14.7 119.3 15.3 84.4 18.0
Specialist SOTAs
Qwen-VL ${}_{\text{Qwen-7B}}$ [5] 1.4B 9.6B 121.4 85.8
CogVLM ${}_{\text{Vicuna-7B}}$ [41] 1.5B 6.5B 148.7 132.6 128.3 94.9
PaLI ${}_{\text{mT5-XXL}}$ [9] 1.6B 17B 149.1 127.0
PaLI-X ${}_{\text{UL2-32B}}$ [8] 2.2B 55B 149.2 126.3
4.2 Evaluation setup
Evaluation dataset. We evaluate EVCap, trained using the COCO training set, across four datasets: its test set, two challenging benchmarks – NoCaps validation set and Flickr30k test set, and a synthetic commonsense-violating dataset – WHOOPS. We adhere follow prior work [15, 41] to use the same images of Karpathy split [22] on COCO test set, NoCaps [2] validation set, and Karpathy split on Flickr30k [30] test set. In addition, WHOOPS [6] is a synthetic image captioning dataset comprising 500 synthetic commonsense-violating images and 2500 paired captions.
Compared methods. We compare EVCap with several SOTAs. According to the trainable parameters size, they can be divided into 1) Heavyweight-training (between 100M to 5B): VinVL [45], AoANet [16], NOC-REK [40], RCA-NOC [13], ViECap [15], InstructBLIP [11], OSCAR [26], BLIP [24], BLIP-2 [25], REVEAL [20]; 2) Lightweight-training (less than 100M): MiniGPT4 [46], SmallCap [35], ClipCap [29]; and also 3) Specialist SOTAs with huge trainable parameters (larger than 5B): Qwen-VL [5], CogVLM [41], PaLI [9], PaLI-X [8]. Among these methods, AoANet, NOC-REK, RCA-NOC, REVEAL, and SmallCap are retrieval-augmented captioning methods.
5 Experimental Results
5.1 Results on in-/out-domain benchmarks
Figure 3: Examples of captions generated by our EVCap and three SOTA methods on COCO test set, NoCaps validation set, and Flickr30k test set. GT refers to the Ground Truth captions. Incorrect objects in captions are highlighted in red, while correct ones are in blue. Our EVCap correctly generates captions across different datasets, showing performance comparable to BLIP-2. In contrast, SmallCap and MiniGPT4 sometimes either miss objects or include incorrect ones in their generated captions.
We assess EVCap against SOTAs on both in-domain and out-domain benchmarks. The COCO test set can be considered as in-domain data as we only train our model on the COCO training set. Out-domain benchmarks are the NoCaps validation set and the Flickr30k test set.
Quantitative results. Tab. 1 details our EVCap’s performance in comparison with SOTA methods. We first evaluate training costs in terms of training data sizes and parameters. Similar to various heavyweight-training models that exclude LLMs and the majority of lightweight-training models, EVCap is trained solely on the COCO training set. It utilizes only 3.97M trainable parameters, positioning it as the second smallest, slightly larger than MiniGPT4 with 3.94M. Among lightweight-training models, our approach outperforms others, achieving the highest scores on all benchmarks. Despite using less training data and nearly identical trainable parameters as MiniGPT4, EVCap significantly surpasses it, with a marked improvement of 10.5, 10.5, and 6.0 in CIDEr scores for each benchmark. When further compared with heavyweight-training models, the performance of EVCap stands out among million-level models, nearly matching InstructBLIP, except in NoCaps. Note that since BLIP-2 does not include Vicuna checkpoints, InstructBLIP performs pre-training with Vicuna using the same procedure as BLIP-2, whereas EVCap does not involve pre-training. Against REVEAL, which also uses external memory, our EVCap utilizes about 1/3000 training data and 1/500 training parameters yet yields comparable results. Moreover, EVCap’s performance is on par with BLIP-2, the top-performing model with 1.2B trainable parameters. This highlights EVCap’s efficiency and effectiveness despite its significantly smaller training cost, thanks to our external visual–name memory. Regarding specialist SOTAs, they use billion-level training data and over 5B trainable parameters, so it is acceptable that they can achieve exceptionally strong performance, surpassing EVCap by nearly 10 on all benchmarks in CIDEr scores.
Qualitative results. Fig. 3 presents a comparison of captions generated by our EVCap and three SOTA models across three benchmarks. The captions of SmallCap are generated by its publicly accessible demo [1]. We generate captions of MiniGPT4 and BLIP-2 using their respective pre-trained models. As a lightweight and retrieval-augmented captioning method, SmallCap struggles to produce accurate captions for given images, primarily because it relies on retrieved captions laden with extraneous information. MiniGPT4, though aligned with the primary content of images, sometimes misses certain objects like “trees” and “headphones”. This oversight stems from its focus on the main objects in images, without integrating additional cues for other objects provided by the retrieved object names. In contrast, the captions generated by our EVCap are comparable to those of BLIP-2.
5.2 Results on commonsense-violating data
To explore our EVCap’s capability in describing contents in open-word settings, we further evaluate it on WHOOPS dataset, which contains commonsense-violating images.
Table 2: Quantitative results on commonsense-violating data – WHOOPS dataset. EVCap (w/ WHOOPS) denotes EVCap using the memory expanded by WHOOPS objects. The results reveal the open-world comprehension ability and expandability of EVCap.
Method B@4 M C S
Only pre-trained models
BLIP  [24] (from [6]) 13 65
BLIP-2 ${}_{\text{FlanT5-XXL}}$ [25] (from [6]) 31 120
BLIP-2 ${}_{\text{FlanT5-XXL}}$ [25] (reproduced) 28 26.7 93.1 17.9
Finetuned models on COCO
MiniGPT4 [46] 24.2 26.7 84.8 18.2
BLIP  [24] 22.9 25.0 79.3 17.1
BLIP-2 ${}_{\text{FlanT5-XL}}$ [25] 25.8 27.0 89.1 18.3
End-to-end trained models on COCO
EVCap 24.1 26.1 85.3 17.7
EVCap (w/ WHOOPS) 24.4 26.1 86.3 17.8
Quantitative results. In Tab. 2, we compare the performance of EVCap, MiniGPT4, BLIP, and BLIP-2 on WHOOPS dataset. This dataset is particularly challenging due to its inclusion of unusual objects [6]. Initially, as an end-to-end trained model, our EVCap exhibits performance similar to MiniGPT4. However, there is a noticeable improvement in the CIDEr score, after the external memory is enriched with 2396 new objects from the WHOOPS dataset, each represented by 5 synthesized images generated using stable diffusion [36]. It highlights the effectiveness of our idea of incorporating an expandable external memory into the captioning model for open-world comprehension.
Qualitative results. Fig. 4 illustrates the captions generated by EVCap, EVCap (w/WHOOPS), and three SOTAs for two images from the WHOOPS dataset. The first image challenges common sense as it unusually pairs “Einstein” with “racing car”. While all SOTAs simply refer to “Einstein” as “a man”, our EVCap and EVCap (w/WHOOPS) correctly identify him. The second image shows another unusual example. Similar to other methods except for BLIP-2, EVCap can not recognize “blue cartoon character” as “Pikachu”, while EVCap (w/WHOOPS) successfully predicts it because of the updated memory. In these two images, SmallCap and MiniGPT4 tend to generate captions with hallucinatory objects, a result of commonsense-violating contents present in the images.
Figure 4: Examples of captions generated by our EVCap, EVCap (w/ WHOOPS), and three SOTAs on WHOOPS dataset. Incorrect objects are highlighted in red, while correct ones are in blue.
5.3 Detailed analysis
Table 3: Ablation study on components prior to the LLM decoder in EVCap. The result of “+ Attentive fusion” demonstrates the substantial impact of the external visual–name memory.
Method COCO test NoCaps val Flickr30k test
C S C S C S
ViT + Q-Former (Baseline) 134.4 23.9 108.8 14.2 76.8 17.3
    + Image query tokens (Baseline+) 134.1 23.8 109.0 14.3 77.3 17.2
    + Attentive fusion (EVCap) 140.1 24.7 119.3 15.3 84.4 18.0
Figure 5: Visualization of the captions generated from ablation study on the NoCaps validation set. We also show the retrieved object names by EVCap, presented in gray. Incorrect objects in captions are highlighted in red, while correct ones are in blue.
Ablation study. We assess the contribution of each component prior to the LLM decoder in EVCap by incrementally integrating the image query tokens and the attentive fusion module into our baseline model. The baseline model comprises a ViT+Q-Former, a linear layer, and a LLM decoder. The quantitative results are shown in Tab. 3. When employing only the baseline model (Baseline), CIDEr scores drop notably by 5.7, 10.5, and 7.6 on COCO, NoCaps, and Flickr30k, respectively. The inclusion of trainable image query tokens (Baseline+) brings a marginal improvement on NoCaps and Flickr30k. However, the performance is significantly enhanced with the addition of attentive fusion (along with the introduction of external memory), indicating the pivotal role of the external visual–name memory in the overall effectiveness of EVCap. This is further corroborated by the qualitative results in Fig. 5, where captions from Baseline and Baseline+ inaccurately include objects like “couch” and “bed”, and Baseline+ overlooks “hand”.
Figure 6: Visualization of the visual features in external memory using t-SNE. For visual features in LVIS dataset’s objects (blue), the related objects fall in the same cluster. After adding more visual features of synthesized images from WHOOPS’ objects, new objects (red) are located at appropriate clusters (zoom-in view).
Exploration for external memory expandability. To demonstrate the scalability of the external memory in EVCap, we visualize the visual features stored in LVIS external memory, and newly synthesized data from objects appearing in the WHOOPS dataset. We employ t-SNE [38] to plot visual features after reducing their dimensions to 2-D (Fig. 6). For clear visualization, we only randomly display 3649 visual features in LVIS memory, and add 479 visual features from WHOOPS objects. Among them, 35 samples are randomly labeled. The result shows a clear clustering of LVIS objects (blue) in the external memory, as well as the successful integration and appropriate localization of new objects from WHOOPS (red) into these clusters. This pattern not only confirms the distinctiveness of visual features already present in the memory but also demonstrates the potential to accurately incorporate and differentiate new objects introduced from updated data. These findings highlight our external memory’s ability to expand and maintain its effectiveness even as new data is incorporated.
Impact of external memory size. We examine the impact of external memory size in Tab. 4. On the one hand, we randomly remove 30%, 60%, and 90% data in the external memory constructed from LVIS objects. The results show the performance gradually degrades on NoCaps as reducing 30% and 90% LVIS. Despite some unexpected increases in certain results on NoCaps (5th row) and Flickr30k (4th - 5th rows), they do not alter the overall downward trend. Similar phenomena are also noted in SmallCap [35], we speculate it is due to data distribution. On the other hand, as we infuse WHOOPS knowledge into LVIS memory, there is a slight improvement on NoCaps (out) and Flickr30k. These observations validate the model’s capability to effectively retrieve object names from an updated memory, enhancing its performance in generating captions.
Table 4: Impact of the external memory size on the performance of EVCap by evaluation under CIDEr scores. Changes in the size of external memory result in changes in performance.
Method NoCaps val Flickr30k
In Near Out Overall Test
LVIS objects (EVCap) 111.7 119.5 116.5 119.3 84.4
    – 30% LVIS 112.0 119.2 115.3 118.8 85.0
    – 60% LVIS 111.4 119.1 116.2 119.0 85.1
    – 90% LVIS 110.6 118.2 115.8 118.3 83.6
    + WHOOPS 110.7 118.9 116.7 119.0 84.9
Figure 7: CIDEr scores after training EVCap with the number of retrieved object names $\mathrm{K}$ from 1 to 20. The results indicate that the performance is relatively optimal when $\mathrm{K}$ is set to be 10.
Impact of the number of retrieved object names. We investigate how the number of retrieved object names $\mathrm{K}$ (Sec. 3.3) affect EVCap in Fig. 7. We train the model with $\mathrm{K}$ from 1 to 20 and evaluate the performance under CIDEr on all three benchmarks. From the results, we can find that the model works worst on the out-domain dataset (NoCaps) when only one retrieved object name is used. As we gradually add more object names, performance fluctuates but improves. This pattern aligns with our intuition that when using one object name, the model will make errors or miss some objects in generated captions due to the incorrect object name. When we add more object names, we increase the error tolerance, and the attentive module in EVCap automatically pays more attention to image-related object names, thus improving results. Furthermore, we observe that setting $\mathrm{K}$ to 10 yields relatively optimal overall performance, validating the choice of $\mathrm{K}=10$ in EVCap.
Analysis with different decoders. To explore the influence of different LLMs decoders on our EVCap, we experiment by substituting Vicuna-13B with GPT2 and Vicuna-7B, as detailed in Tab. 5. With GPT2 as the decoder, EVCap still markedly surpasses other GPT2-based models, achieving impressive gains of 11.3 and 10.0 under CIDEr on COCO and Flickr30k, compared to SmallCap. When employing Vicuna-7B, the comparison of performance trends mirrors those observed with Vicuna-13B, further attesting to the robustness and adaptability of EVCap across different LLM decoders. Notably, both SmallCap, which retrieves captions, and our GPT2-based EVCap, which retrieves object names, use the same GPT2 decoder. Therefore, their comparison also underscores the effectiveness of our method’s object name retrieval and attentive fusion strategy.
Table 5: Analysis with different LLM decoders including GPT2, Vicuna-7B, and Vicuna-13B. The results reveal EVCap is effective when applying it in different LLM decoders.
Method LLM COCO test NoCaps val Flickr30k test
C S C S C S
SmallCap [35] GPT2 119.7 21.3 60.6
ViECap [15] GPT2 92.9 18.2 66.2 9.5 47.9 13.6
EVCap GPT2 131.0 23.2 97.6 13.3 70.6 16.1
MiniGPT4 [46] Vicuna-7B 119.4 23.5 108.7 15.7 73.9 17.2
InstructBLIP [11] Vicuna-7B 123.1 82.4
EVCap Vicuna-7B 139.0 24.7 116.8 15.3 82.7 18.0
MiniGPT4 [46] Vicuna-13B 129.6 23.4 108.8 15.1 78.4 16.9
InstructBLIP [11] Vicuna-13B 121.9 82.8
EVCap Vicuna-13B 140.1 24.7 119.3 15.3 84.4 18.0
Limitations. First, EVCap cannot retrieve all objects that appear in the given image, leading to incomplete image descriptions as the second example in Fig. 4. We will investigate integrating object detection with image captioning to enhance completeness. Second, our focus on object representation restricts consideration of other crucial captioning elements, affecting overall performance. Similar to all models trained on the COCO dataset, our EVCap inherits its captioning style, which limits its ability to generate varied styles. This limitation is reflected in our relatively modest performance improvements in Tab. 2, compared to MiniGPT4. We will overcome this limitation by exploring methodologies that encourage style diversity in the future.
6 Conclusion
We further advance image captioning in real-world scenarios by introducing EVCap, a novel image captioning model with object names retrieved from an external visual–name memory. The external memory is easily expandable, allowing for effortless updates with new object visuals and names. EVCap stands out for its efficiency, comprising merely 3.97M trainable parameters, yet delivering robust performance. We extensively compare EVCap with SOTAs on various benchmarks and commonsense-violating data, demonstrating its significant superiority in performance.
Supplementary Material for
EVCap: Retrieval-Augmented Image Captioning with External Visual–Name Memory for Open-World Comprehension
Figure A: Samples of images and corresponding names in the external visual–name memory constructed from LVIS’s objects.
This supplementary material complements our paper with the following sections: First, we delve into the implementation specifics of our EVCap, which were not covered in the main paper (see Sec. A). Second, we offer an expanded discussion on the external visual-name memory, as utilized in the main paper (see Sec. B). Finally, we present additional results to evaluate the effectiveness of EVCap (see Sec. C).
A Implementation Details
Our method is based on Pytorch and is trained within one epoch with a batch size of 24 using mixed precisions. We optimize the model using AdamW, setting the weight decay at 0.05, and using $\beta_{1}$ and $\beta_{2}$ values of 0.9 and 0.99, respectively. A cosine learning rate (LR) decay strategy is adopted, starting with an initial LR of 1e-4. The model undergoes 5000 linear warm-up steps, beginning with a start LR of 1e-6. During the evaluation phase, we use a beam search strategy with a beam size of 5 to generate captions.
Figure B: Samples of images and corresponding names in the added WHOOPS’s objects.
B External Visual–Name Memory
B.1 LVIS memory
As stated in Sec. 3.2 of the main paper, we utilize 1203 objects from the LVIS dataset. For each of these objects, we randomly select between one and ten images from LVIS. Additionally, we enrich our data by incorporating five synthetic images for each object, created using stable diffusion. We show two samples of this external visual-name memory, constructed using objects from LVIS in Fig. A.
B.2 WHOOPS memory
To illustrate the scalability of the external memory in EVCap, we expand it by integrating WHOOPS knowledge into the original external visual–name memory in Sec. 5.2 and Sec. 5.3 of the main paper. Specifically, we focus on objects that are mentioned in the answers of VQA annotations in the WHOOPS dataset because of their conciseness and emphasis on key objects. For each of these objects, we produce five synthetic images employing stable diffusion. Two examples from this augmented memory, featuring newly added object images and their corresponding names, are presented in Fig. B.
C Additional Results
[…]
Method In Near Out Overall
Heavyweight-training models
VinVL [45] 93.8 89.0 66.1 85.5
NOC-REK* [40] 100.0 95.7 77.4 93.0
OSCAR [26] 81.3 79.6 73.6 78.8
Lightweight-training models
SmallCap* ${}_{\text{GPT2}}$ [35] 87.9 84.6 84.4 85.0
EVCap* ${}_{\text{Vicuna-13B}}$ 114.9 117.0 117.1 116.8
Specialist SOTAs
CogVLM ${}_{\text{Vicuna-7B}}$ [5] 128.0 126.4
PaLI ${}_{\text{mT5-XXL}}$ [9] 124.4
PaLI-X ${}_{\text{UL2-32B}}$ [8] 124.3
${}^{a}$https://eval.ai/web/challenges/challenge-page/355/leaderboard/1011
C.1 Experiments on NoCaps test set
We additionally assess our EVCap against SOTAs on the NoCaps test set, since we notice several other methods have also benchmarked their performance on this dataset. Note that, NoCaps test set does not have publicly accessible ground truth annotations. To obtain evaluation scores, we submitted our results to the NoCaps leaderboard.
Quantitative results. Tab. A presents the quantitative results of our EVCap on the NoCaps test set. Our method outperforms all other SOTA models, both in heavyweight-training and lightweight-training categories, that have reported results on this dataset. Additionally, as a lightweight method, our approach achieves the 8th rank on the NoCaps leaderboard, only surpassed by specialized SOTAs such as CogVLM, which holds the 1st rank.
Qualitative results. Fig. D shows the captions generated by our EVCap alongside those from three SOTA methods on the NoCaps test set. It also includes the object names retrieved by EVCap and the captions retrieved by SmallCap. Consistent with the findings in the main paper, SmallCap tends to generate hallucinatory objects that are absent in the input images, such as “tie” and “mouse”. The same hallucinatory object “mouse” is also found in the retrieved captions, indicating that SmallCap’s diminished performance is largely due to its reliance on retrieved captions containing irrelevant information. In comparison, our EVCap demonstrates a performance on par with BLIP-2.
Figure C: CIDEr scores after training EVCap with the number of object name query tokens $P$ from 2 to 10. The results indicate that the performance is relatively optimal when $P$ is set to be 8.
Figure D: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
Table B: Comparison against SOTA methods on training time and used GPUs. * denotes using a memory bank.
Method Training time GPUs
Heavyweight-training models
VinVL [45]
AoANet+MA* [16]
NOC-REK* [40] 8d 2 RTX3090
RCA-NOC*[13] 1d 8 A100
ViECap ${}_{\text{GPT2}}$[15]
InstructBLIP ${}_{\text{Vicuna-13B}}$[11] 1.5d 16 A100
OSCAR [26] 74h 1 V100
BLIP [24] 2 16-GPU nodes
BLIP-2 ${}_{\text{FlanT5-XL}}$ [25] $\sim$9d 16 A100
REVEAL* ${}_{\text{T5}}$ [20] 5d 256 CloudTPUv4 chips
Lightweight-training models
MiniGPT4 ${}_{\text{Vicuna-13B}}$ [46] 10h 4 A100
SmallCap* ${}_{\text{GPT2}}$ [35] 8h 1 A100
ClipCap ${}_{\text{GPT2}}$ [29] 6h 1 GTX1080
EVCap* ${}_{\text{Vicuna-13B}}$ 3h 4 A6000
Specialist SOTAs
Qwen-VL ${}_{\text{Qwen-7B}}$ [5]
CogVLM ${}_{\text{Vicuna-7B}}$ [41] 1d 4096 A100
PaLI ${}_{\text{mT5-XXL}}$ [9]
PaLI-X ${}_{\text{UL2-32B}}$ [8]
C.2 Further analysis
Comparison on training time and used GPUs. Tab. B compares training time and the used GPUs of our EVCap with various SOTA models. Due to the diversity of GPUs employed across different models, drawing a direct comparison is challenging. Nevertheless, it’s evident that the training time for our EVCap is comparatively shorter than most models.
Number of object name query tokens. We explore the impact of varying the number of retrieved object names $P$ (Sec. 3.4 of the main paper) on EVCap in Fig. C. We train the model using different values of $P$, ranging from 2 to 10, and evaluate the performance under CIDEr on all three benchmarks. The results suggest that setting $P=8$ offers relatively optimal results.
Figure E: Examples of captions generated by our EVCap, and three SOTAs on the COCO test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
Figure F: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps validation set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
C.3 More qualitative examples.
More qualitative examples on the COCO test set, NoCaps validation set, and Flickr30k test set are shown in Fig. E, Fig. F, and Fig. G, respectively.
Figure G: Examples of captions generated by our EVCap, and three SOTAs on the Flickr30k test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.

### 07 Apr 2024 v2

@@
 Large language models (LLMs)-based image captioning has the capability of describing objects not explicitly observed in training data; yet novel objects occur frequently, necessitating the requirement of sustaining up-to-date object knowledge for open-world comprehension.
-Instead of relying on large amounts of data and scaling up network parameters, we introduce a highly effective retrieval-augmented image captioning method that prompts LLMs with object names retrieved from External Visual–name memory (EVCap).
-We build ever-changing object knowledge memory using objects’ visuals and names, enabling us to (i) update the memory at a minimal cost and (ii) effortlessly augment LLMs with retrieved object names utilizing a lightweight and fast-to-train model.
-Our model, which was trained only on the COCO dataset, can be adapted to out-domain data without additional fine-tuning or retraining.
-Our comprehensive experiments conducted on various benchmarks and synthetic commonsense-violating data demonstrate that EVCap, comprising solely 3.97M trainable parameters, exhibits superior performance compared to other methods of equivalent model size scale.
-Notably, it achieves competitive performance against specialist SOTAs with an enormous number of parameters.
-Our code is available at https://jiaxuan-li.github.io/EVCap.
+Instead of relying on large amounts of data and/or scaling up network parameters, we introduce a highly effective retrieval-augmented image captioning method that prompts LLMs with object names retrieved from External Visual–name memory (EVCap).
+We build ever-changing object knowledge memory using objects’ visuals and names, enabling us to (i) update the memory at a minimal cost and (ii) effortlessly augment LLMs with retrieved object names by utilizing a lightweight and fast-to-train model.
+Our model, which was trained only on the COCO dataset, can adapt to out-of-domain without requiring additional fine-tuning or re-training.
+Our experiments conducted on benchmarks and synthetic commonsense-violating data show that EVCap, with only 3.97M trainable parameters, exhibits superior performance compared to other methods based on frozen pre-trained LLMs.
+Its performance is also competitive to specialist SOTAs that require extensive training.
+*Equal contributions.
+Code is available at https://jiaxuan-li.github.io/EVCap.
 1 Introduction
 Advanced image captioning based on large language models (LLMs) [3, 9, 25, 8] has focused on the approach using big-scale models trained on ever-increasingly large-scale datasets, which is no longer viable.
-This is because the computational cost to train the models increases exponentially and, more importantly, updating training data is almost impossible that keep pace with the growth of novel objects in our daily lives.
+This is because the computational cost to train the models increases exponentially and, more importantly, updating training data is almost impossible to keep pace with the growth of novel objects in our daily lives.
 Sustaining ever-changing object knowledge with a reasonable cost is a pressing concern in LLMs-based models to truly unlock open-world comprehension.
 Figure 1: Overall comparison of our EVCap and SOTAs.
-(Upper) Comparison of the number of trainable parameters, CIDEr score on COCO and NoCaps datasets.
+(Upper) Generated captions by SmallCap, BLIP-2, and our EVCap for a commonsense-violating image from the WHOOPS dataset.
+$\times$ and $\checkmark$ indicate incorrect and correct predictions, respectively.
+Incorrect objects in captions are highlighted in red, while correct ones are in blue.
+SmallCap and BLIP-2 give incorrect predictions for “ice skates" and “wood floor", respectively, while our EVCap utilizes an external visual–name memory to enhance attention to objects within the image, leading to superior performance for image captioning.
+(Lower) Comparison of the number of trainable parameters, CIDEr score on COCO and NoCaps datasets.
 The size of each circle reflects the log number of trainable parameters.
 EVCap (3.97M) has less trainable parameters than others while achieving comparable results with SOTAs at scale.
-(Lower) Generated captions by SmallCap, BLIP-2, and our EVCap for a commonsense-violating image from the WHOOPS dataset.
-$\times$ and $\checkmark$ indicate incorrect and correct predictions, respectively.
-Incorrect objects in captions are highlighted in red, while correct ones are in blue.
-SmallCap and BLIP-2 give incorrect predictions for “ice skates” and “wood floor”, respectively, while our EVCap utilizes an external visual–name memory to enhance attention to objects within the image, leading to superior performance for image captioning.
 Retrieval-augmented image captioning [35, 20] is emerging as an alternative since it considerably reduces training costs in both time and data while producing encouraging results.
 Nonetheless, with their huge datastore, it is obvious that LLMs would imitate the given texts, limiting their ability to describe open-world objects properly.
-For instance, SmallCap [35] considers the words “skateboard” and “wooden floor” to be a pair regardless of visual appearances containing a commonsense-violating pair of “ice skates” and “wood floor” (Fig.
-1, lower).
+For instance, SmallCap [35] considers the words “skateboard" and “wooden floor" to be a pair regardless of visual appearances containing a commonsense-violating pair of “ice skates" and “wood floor" (Fig.
+1, upper).
 Additionally, prompting the LLMs given a lot of retrieved texts becomes cumbersome, requiring more trainable parameters.
 Fig.
-1 (upper) shows that the CIDEr scores obtained by a lightweight SmallCap [35] with 43M trainable parameters are far away from those obtained by a heavy REVEAL [20] with 2.1B trainable parameters.
+1 (lower) shows that the CIDEr scores obtained by a lightweight SmallCap [35] with 43M trainable parameters are far away from those obtained by a heavy REVEAL [20] with 2.1B trainable parameters.
 Beyond that, due to the frequent occurrence of new objects, access to their sample texts is not always feasible, making the memory utilized in [35, 20] difficult to grow.
@@
 As a result, EVCap amounts to only 3.97M trainable parameters.
-Once trained, the model can be adapted to new domains and large-scale data without further fine-tuning or retraining.
+Once trained, the model can be adapted to new domains and large-scale data without further fine-tuning or re-training.
 Our contributions are as follows:
   • We provide an extensible external visual–name memory with minimal but useful object information, which enables LLMs-based models to comprehend the open world.
-  • We present a remarkably lightweight and highly efficacious retrieval-augmented image captioning EVCap with 3.97M trainable parameters.
+  • We present a highly efficacious retrieval-augmented image captioning EVCap with 3.97M trainable parameters.
 On in-/out-domain benchmarks and synthetic commonsense-violating dataset, EVCap trained solely on COCO dataset competes with other lightweight methods by a margin while being on par with other specialist SOTAs.
@@
 Novel object captioning is a branch of image captioning that describes images containing objects that were not seen during training.
-Non-LLMs-based methods explore more objects by learning from unpaired image-sentence sources (DCC [19], NOC [39]) or relied on novel object detectors to recognize novel concepts (NBT [28], OSCAR [26] and VinVL [45]).
+Non-LLMs-based methods explore more objects by learning from unpaired image-sentence sources (DCC [19], NOC [39]) or rely on novel object detectors to recognize novel concepts (NBT [28], OSCAR [26] and VinVL [45]).
 LLMs-based methods such as ViECap [15] leverage the pre-trained CLIP [32] to obtain object entities.
@@
 EXTRA [34] and Re-ViLM [43] exploit the similarity of the input image and vision candidates to retrieve captions.
-Unlike previous methods, our external memory contains visual–name pairs to avoid redundant information in the external captions/documents.
+Different from the previous methods, our external memory contains visual–name pairs to avoid redundant information in the external captions/documents.
 In addition, we use an attentive fusion module to mitigate the effects of irrelevant retrieved object names on caption generation.
@@
 We aim to build a retrieval-augmented LLMs-based image captioning model with a sufficiently small yet informative external memory.
-It involves two challenges: (1) constructing an expandable external memory, and (2) building an effective LLMs-based model using retrieved object names.
+It involves two challenges: (1) constructing an expandable external memory and (2) building an effective LLMs-based model using retrieved object names.
 As discussed above, challenge (1) can be resolved by utilizing the visual appearance of objects.
@@
 With the capability to collect images from the internet, EVCap can be easily expanded to include novel objects from the real world effortlessly.
-We base our method on a frozen pre-trained vision model and LLM with several trainable layers (Fig.
+We base our method on the frozen pre-trained vision model and LLM with several trainable layers (Fig.
 2), giving in a model that is cheap to train.
@@
 These selected values $v_{j}^{\mathrm{best}}$ are redefined as the new notation $v_{l}$ in the retrieved top-K object names for the input image, which can be summarized as follows:
-$\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}=\arg\max_{\mathbf{k}^{i}}\operatorname{SIM}\left(\mathbf{q}_{j},\mathbf{k}^{i}\right),\\ \vspace*{-0.75\baselineskip}$
+$\{\mathbf{k}_{j}^{\mathrm{best}},v_{j}^{\mathrm{best}}\}=\arg\max_{\mathbf{k}^{i}}\operatorname{SIM}\left(\mathbf{q}_{j},\mathbf{k}^{i}\right),\$
 $j=\arg\max_{j}\operatorname{SIM}(\mathbf{q}_{j},\mathbf{k}_{j}^{\mathrm{best}}),v_{l}\leftarrow v_{j}^{\mathrm{best}},$
@@
 The sequence $\mathcal{S}$ and visual features $\mathcal{Q}$ are fed into a customized Q-Former $\mathcal{F(\cdot)}$, which is constructed from the frozen pre-trained Q-Former as we used in vision encoder $\mathcal{E}$.
-Nonetheless, in order to enable object names to get attention from visual features, we switch the image embedding port and the text instruction port (see [25] for architecture detail).
+Nonetheless, in order to enable object names to get attention from visual features, we switch the image embedding port and the text instruction port (see the supplement for architecture detail).
 Like in the image encoding process in Sec.
@@
 For all experiments, we exclusively train EVCap using the training set of COCO dataset [27], consisting of 82k images and 5 captions per images.
-The entire training process takes about 3 hours on 4 A6000 GPUs, using mixed precisions (more details in the supplementary).
+The entire training process takes about 3 hours on 4 A6000 GPUs, using mixed precisions (more details in the supplement).
 Table 1: Quantitative comparison against SOTA methods on three common image captioning benchmarks.
@@
 Our EVCap correctly generates captions across different datasets, showing performance comparable to BLIP-2.
-In contrast, SmallCap and MiniGPT4 sometimes either miss objects or include incorrect ones in their generated captions.
 We assess EVCap against SOTAs on both in-domain and out-domain benchmarks.
@@
 As a lightweight and retrieval-augmented captioning method, SmallCap struggles to produce accurate captions for given images, primarily because it relies on retrieved captions laden with extraneous information.
-MiniGPT4, though aligned with the primary content of images, sometimes misses certain objects like “trees” and “headphones”.
+MiniGPT4, though aligned with the primary content of images, sometimes misses certain objects like “trees" and “headphones".
 This oversight stems from its focus on the main objects in images, without integrating additional cues for other objects provided by the retrieved object names.
@@
 To explore our EVCap’s capability in describing contents in open-word settings, we further evaluate it on WHOOPS dataset, which contains commonsense-violating images.
+Figure 4: Examples of captions generated by our EVCap, EVCap (w/ WHOOPS), and three SOTAs on WHOOPS dataset. Incorrect objects are highlighted in red, while correct ones are in blue.
 Table 2: Quantitative results on commonsense-violating data – WHOOPS dataset.
@@
 Fig.
-4 illustrates the captions generated by EVCap, EVCap (w/WHOOPS), and three SOTAs for two images from the WHOOPS dataset.
-The first image challenges common sense as it unusually pairs “Einstein” with “racing car”.
-While all SOTAs simply refer to “Einstein” as “a man”, our EVCap and EVCap (w/WHOOPS) correctly identify him.
-The second image shows another unusual example.
-Similar to other methods except for BLIP-2, EVCap can not recognize “blue cartoon character” as “Pikachu”, while EVCap (w/WHOOPS) successfully predicts it because of the updated memory.
-In these two images, SmallCap and MiniGPT4 tend to generate captions with hallucinatory objects, a result of commonsense-violating contents present in the images.
-Figure 4: Examples of captions generated by our EVCap, EVCap (w/ WHOOPS), and three SOTAs on WHOOPS dataset. Incorrect objects are highlighted in red, while correct ones are in blue.
+4 illustrates the captions generated by EVCap, EVCap (w/ WHOOPS), and three SOTAs for one image from the WHOOPS dataset.
+Similar to other methods except for BLIP-2, EVCap cannot recognize “blue cartoon character" as “Pikachu", while EVCap (w/WHOOPS) successfully predicts it because of the updated memory.
+SmallCap and MiniGPT4 tend to generate captions with hallucinatory objects, a result of commonsense-violating contents present in the images.
 5.3 Detailed analysis
-Table 3: Ablation study on components prior to the LLM decoder in EVCap. The result of “+ Attentive fusion” demonstrates the substantial impact of the external visual–name memory.
+Table 3: Ablation study on components prior to the LLM decoder in EVCap. The result of “+ Attentive fusion" demonstrates the substantial impact of the external visual–name memory.
 Method COCO test NoCaps val Flickr30k test
@@
 This is further corroborated by the qualitative results in Fig.
-5, where captions from Baseline and Baseline+ inaccurately include objects like “couch” and “bed”, and Baseline+ overlooks “hand”.
+5, where captions from Baseline and Baseline+ inaccurately include objects like “couch" and “bed", and Baseline+ overlooks “hand".
 Figure 6: Visualization of the visual features in external memory using t-SNE.
@@
     + WHOOPS 110.7 118.9 116.7 119.0 84.9
-Figure 7: CIDEr scores after training EVCap with the number of retrieved object names $\mathrm{K}$ from 1 to 20.
+Figure 7: CIDEr scores after training EVCap with the number of retrieved object names $\mathrm{K}$ from 0 to 20.
 The results indicate that the performance is relatively optimal when $\mathrm{K}$ is set to be 10.
@@
 7.
-We train the model with $\mathrm{K}$ from 1 to 20 and evaluate the performance under CIDEr on all three benchmarks.
-From the results, we can find that the model works worst on the out-domain dataset (NoCaps) when only one retrieved object name is used.
-As we gradually add more object names, performance fluctuates but improves.
-This pattern aligns with our intuition that when using one object name, the model will make errors or miss some objects in generated captions due to the incorrect object name.
-When we add more object names, we increase the error tolerance, and the attentive module in EVCap automatically pays more attention to image-related object names, thus improving results.
+We train the model with $\mathrm{K}$ from 0 to 20 and evaluate the performance under CIDEr on all three benchmarks.
+From the results, we can find that the model works worst on the out-domain dataset (NoCaps) with zero object names.
+It confirms that the performance boost from Baseline+ to EVCap (Tab.
+3) is primarily attributed to the retrieval-augmented mechanism, but not the customized Q-Former itself.
+With more object names, performance fluctuates but improves.
 Furthermore, we observe that setting $\mathrm{K}$ to 10 yields relatively optimal overall performance, validating the choice of $\mathrm{K}=10$ in EVCap.
@@
 Limitations.
-First, EVCap cannot retrieve all objects that appear in the given image, leading to incomplete image descriptions as the second example in Fig.
-4.
+First, EVCap cannot retrieve all objects that appear in the given image due to the memory coverage limits, leading to incomplete image descriptions (Fig.
+4).
 We will investigate integrating object detection with image captioning to enhance completeness.
 Second, our focus on object representation restricts consideration of other crucial captioning elements, affecting overall performance.
-Similar to all models trained on the COCO dataset, our EVCap inherits its captioning style, which limits its ability to generate varied styles.
-This limitation is reflected in our relatively modest performance improvements in Tab.
+Similar to all models trained with COCO dataset, EVCap has limitations in generating varied styles, which is reflected in our relatively modest performance improvements in Tab.
 2, compared to MiniGPT4.
-We will overcome this limitation by exploring methodologies that encourage style diversity in the future.
+We will overcome it by exploring methodologies that encourage style diversity in the future.
 6 Conclusion
@@
 The external memory is easily expandable, allowing for effortless updates with new object visuals and names.
-EVCap stands out for its efficiency, comprising merely 3.97M trainable parameters, yet delivering robust performance.
 We extensively compare EVCap with SOTAs on various benchmarks and commonsense-violating data, demonstrating its significant superiority in performance.
+Acknowledgements.
+This work was supported by JSPS/MEXT KAKENHI Grant Numbers JP22H05015, JP23H03449, JP23KJ0404, and JP22K17947, and the commissioned research (No.
+225) by the National Institute of Information and Communications Technology (NICT), Japan, and the Institute of AI and Beyond of the University of Tokyo, and ROIS NII Open Collaborative Research 2023-23FC01, 2024-24S1201.
 Supplementary Material for
 EVCap: Retrieval-Augmented Image Captioning with External Visual–Name Memory for Open-World Comprehension
-Figure A: Samples of images and corresponding names in the external visual–name memory constructed from LVIS’s objects.
 This supplementary material complements our paper with the following sections: First, we delve into the implementation specifics of our EVCap, which were not covered in the main paper (see Sec.
@@
 A Implementation Details
+Customized Q-Former.
+Fig.
+A depicts customized Q-Former, where the image embedding port (CLIP visual feature port) now receives retrieved object names $\mathcal{S}$, while the text port receives visual features $\mathcal{Q}$.
+Figure A: Customized Q-Former and original Q-Former [25].
+Implementation.
 Our method is based on Pytorch and is trained within one epoch with a batch size of 24 using mixed precisions.
@@
 During the evaluation phase, we use a beam search strategy with a beam size of 5 to generate captions.
-Figure B: Samples of images and corresponding names in the added WHOOPS’s objects.
+Figure B: Samples of images and corresponding names in the external visual–name memory constructed from LVIS’s objects.
+Figure C: Samples of images and corresponding names in the added WHOOPS’s objects.
 B External Visual–Name Memory
@@
 We show two samples of this external visual-name memory, constructed using objects from LVIS in Fig.
-A.
+B.
 B.2 WHOOPS memory
@@
 Two examples from this augmented memory, featuring newly added object images and their corresponding names, are presented in Fig.
-B.
+C.
 C Additional Results
@@
 Fig.
-D shows the captions generated by our EVCap alongside those from three SOTA methods on the NoCaps test set.
+E shows the captions generated by our EVCap alongside those from three SOTA methods on the NoCaps test set.
 It also includes the object names retrieved by EVCap and the captions retrieved by SmallCap.
-Consistent with the findings in the main paper, SmallCap tends to generate hallucinatory objects that are absent in the input images, such as “tie” and “mouse”.
-The same hallucinatory object “mouse” is also found in the retrieved captions, indicating that SmallCap’s diminished performance is largely due to its reliance on retrieved captions containing irrelevant information.
+Consistent with the findings in the main paper, SmallCap tends to generate hallucinatory objects that are absent in the input images, such as “tie" and “mouse".
+The same hallucinatory object “mouse" is also found in the retrieved captions, indicating that SmallCap’s diminished performance is largely due to its reliance on retrieved captions containing irrelevant information.
 In comparison, our EVCap demonstrates a performance on par with BLIP-2.
-Figure C: CIDEr scores after training EVCap with the number of object name query tokens $P$ from 2 to 10. The results indicate that the performance is relatively optimal when $P$ is set to be 8.
-Figure D: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps test set.
+Figure D: CIDEr scores after training EVCap with the number of object name query tokens $P$ from 2 to 10. The results indicate that the performance is relatively optimal when $P$ is set to be 8.
+Figure E: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps test set.
 We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
@@
 3.4 of the main paper) on EVCap in Fig.
-C.
+D.
 We train the model using different values of $P$, ranging from 2 to 10, and evaluate the performance under CIDEr on all three benchmarks.
 The results suggest that setting $P=8$ offers relatively optimal results.
-Figure E: Examples of captions generated by our EVCap, and three SOTAs on the COCO test set.
+Figure F: Examples of captions generated by our EVCap, and three SOTAs on the COCO test set.
 We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
-Figure F: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps validation set.
+Figure G: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps validation set.
 We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
 C.3 More qualitative examples.
-More qualitative examples on the COCO test set, NoCaps validation set, and Flickr30k test set are shown in Fig. E, Fig. F, and Fig. G, respectively.
-Figure G: Examples of captions generated by our EVCap, and three SOTAs on the Flickr30k test set.
+More qualitative examples on the COCO test set, NoCaps validation set, and Flickr30k test set are shown in Fig. F, Fig. G, and Fig. H, respectively.
+Figure H: Examples of captions generated by our EVCap, and three SOTAs on the Flickr30k test set.
 We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.